The difference of two numbers is 14 . The larger number is 3 more than twice the smaller number.Find the numbers.

Answer by amoresroy(361)

The difference of two numbers is 14 . The larger number is 3 more than twice the smaller number.Find the numbers.
Let x = the smaller number
y = the larger number
1st equation:
y - x = 14
2nd equation:
y = 2x + 3
Substitute y = 2x + 3 in equation 1 to get
(2x+3)-x = 14
x = 14-3
x = 11
Substitute x = 11 in equation 2 to get
y = 2(11) + 3
y = 25
The numbers are 11 and 25.